# Boundary-supported radial layering in Hoag-like ring galaxies

### 14 May 2026 v1

@@ Abstract
Hoag-like galaxies exhibit an old compact central component, a depleted gap, and a detached outer ring. We identify a boundary-supported radial-layering mechanism in a shell-deformed Kepler control model. A compact inner boundary supplies the core state, while a localized shell, interpreted as the reduced imprint of externally supplied material settled near a finite circularization radius, needs to create only an internal maximum and a subsequent outer minimum. These act as the gap barrier and ring-supporting well. The onset of this structure is organized by a saddle-node threshold of the critical-point equation. In a $10^{4}$-point Monte Carlo scan, shell-localized boundary-supported candidates occupy finite parameter volume, and none of the localized candidates contains an ordered interior minimum–maximum–minimum subsequence. The same branch gives a scale-free gap-to-ring interval consistent with Hoag’s Object, UGC 4599, and PGC 1000714, but not with the environmentally processed comparison object JO171.
@@
galaxies: peculiar – galaxies: structure – galaxies: evolution – galaxies: kinematics and dynamics
@@
Hoag-like ring galaxies are distinguished not simply by the presence of a ring, but by a clean radial separation: a compact central body, an underdense gap, and a detached outer annulus. This structure is difficult to reduce to the standard ring-formation channels. Nearly head-on collisions generate expanding density waves (Lynds & Toomre, 1976; Appleton & Struck-Marcell, 1996), while bar or oval perturbations organize gas near resonances in disk galaxies (Buta & Combes, 1996). The cleanest Hoag-type systems, however, often show spheroidal or nearly round cores, weak or absent bar signatures, and evidence for externally supplied gas or extended H i reservoirs (Schweizer et al., 1987; Finkelman et al., 2011; Finkelman & Brosch, 2011; Pakdil et al., 2017; Moretti et al., 2018).
@@
be a dimensionless radius. The control potential is
@@
The Kepler term represents the compact-core control limit. The Gaussian term is an effective reduced deformation, not a Poisson-self-consistent gas potential; it represents the radial imprint of externally supplied material that has dissipated, partially circularized, and accumulated near a finite radius. Thus $x_{0}$, $\Delta$, and $\epsilon$ describe the settling radius, radial width, and control strength of the shell.
@@
$ds^{2}=-e^{2\Phi(x)}dt^{2}+e^{-2\Phi(x)}dx^{2}+x^{2}d\varphi^{2},$ (6)
@@
$\frac{d\Phi}{dx}=\frac{\ell^{2}}{x(x^{2}+\ell^{2})}.$ (9)
@@
$\frac{d^{2}x_{i}}{d\tau^{2}}+\Gamma\frac{dx_{i}}{d\tau}=-\partial_{x}U_{\rm eff}(x_{i};\ell_{i},\tau),$ (12)
@@
$E_{r,i}=\frac{1}{2}\left(\frac{dx_{i}}{d\tau}\right)^{2}+U_{\rm eff}(x_{i};\ell_{i},\tau)$ (13)
@@
$\frac{dE_{r,i}}{d\tau}=-\Gamma\left(\frac{dx_{i}}{d\tau}\right)^{2}+\partial_{\tau}U_{\rm eff}.$ (14)
@@
Alt text: Schematic radial profile showing a compact inner boundary, a shell-centered maximum labelled as the gap barrier, and an outer minimum labelled as the ring-supporting well.
@@
with log-uniform sampling for $\mu$, $\ell$, and $\Delta$, uniform sampling for $x_{0}$ and $\epsilon$, and $\Delta/x_{0}\leq 0.5$. For each realization all internal critical points of $W$ were located and classified.
@@
Alt text: Scatter plot of accepted Monte Carlo candidates in shell-localization coordinates, concentrated near gap-to-shell ratio one with ring-to-shell ratio greater than one.
@@
respectively (Schweizer et al., 1987; Finkelman et al., 2011; Finkelman & Brosch, 2011; Pakdil et al., 2017). All three fall within the central model interval. JO171, whose ring has been interpreted as externally formed and is now being reshaped by ram-pressure stripping in a cluster environment, has
@@
Alt text: Probability-density plot of the model gap-to-ring ratio with observed markers for Hoag’s Object, UGC 4599, PGC 1000714, and JO171; the first three overlap the model interval and JO171 lies below it.
@@ Acknowledgements
We thank the authors of the observational studies cited here for making Hoag-like systems a quantitatively testable morphology class. We also thank the developers of the open-source Python and tools used in this work.
@@
The observational quantities used in this work are taken from the cited literature. The Monte Carlo catalogues and plotting scripts underlying this article will be shared by the corresponding authors upon reasonable request.

### 15 May 2026 v2

@@ Abstract
Clean Hoag-like ring galaxies are often characterized by an old compact central component, a depleted gap, and a detached outer ring. We identify a boundary-supported radial-layering mechanism in a shell-deformed Kepler control model. A compact inner boundary supplies the core state, while a localized effective shell deformation, interpreted as the reduced imprint of externally supplied material settled near a finite circularization radius, needs to create only an internal maximum and a subsequent outer minimum. These act as the gap barrier and ring-supporting well. The onset of this structure is organized by a saddle-node threshold of the critical-point equation. In a $10^{4}$-point Monte Carlo scan, shell-localized boundary-supported candidates occupy finite parameter volume under the adopted priors, and none of the localized candidates contains an ordered interior minimum–maximum–minimum subsequence. The same branch gives a scale-free gap-to-ring interval overlapping representative ratios for Hoag’s Object, UGC 4599, and PGC 1000714, but not for the environmentally processed comparison object JO171.
@@
galaxies: peculiar – galaxies: structure – galaxies: kinematics and dynamics – methods: analytical – methods: numerical
@@
Hoag-like ring galaxies are distinguished not simply by the presence of a ring, but by a clean radial separation: a compact central body, an underdense gap, and a detached outer annulus. This structure is difficult to reduce to the standard ring-formation channels. Nearly head-on collisions generate expanding density waves (Lynds and Toomre, 1976; Appleton and Struck-Marcell, 1996), while bar or oval perturbations organize gas near resonances in disk galaxies (Buta and Combes, 1996). The cleanest Hoag-type systems, however, often show spheroidal or nearly round cores, weak or absent bar signatures, and evidence for externally supplied gas or extended H i reservoirs (Schweizer et al., 1987; Finkelman et al., 2011; Finkelman and Brosch, 2011; Pakdil et al., 2017; Moretti et al., 2018).
@@
be a dimensionless radius, where $r$ is the physical galactocentric radius and $R_{\rm s}$ is the radial scale used to nondimensionalize the model. In the observational comparison below, $R_{\rm s}$ is fixed by matching the model ring-supporting radius to the observed ring radius, $R_{\rm s}=R_{\rm ring}^{\rm obs}/x_{\rm r}$. The control potential is
@@
Here $x_{0}$, $\Delta$, and $\epsilon$ describe the settling radius, radial width, and control strength of the shell. The Kepler term represents the compact-core control limit. The Gaussian term is an effective reduced deformation, not a Poisson-self-consistent gas potential; it represents the radial imprint of externally supplied material that has dissipated, partially circularized, and accumulated near a finite radius.
@@
$\mathrm{d}s^{2}=-e^{2\Phi(x)}\mathrm{d}t^{2}+e^{-2\Phi(x)}\mathrm{d}x^{2}+x^{2}\mathrm{d}\varphi^{2},$ (6)
@@
$\frac{\mathrm{d}\Phi}{\mathrm{d}x}=\frac{\ell^{2}}{x(x^{2}+\ell^{2})}.$ (9)
@@
$\frac{\mathrm{d}^{2}x_{i}}{\mathrm{d}\tau^{2}}+\Gamma\frac{\mathrm{d}x_{i}}{\mathrm{d}\tau}=-\partial_{x}U_{\rm eff}(x_{i};\ell_{i},\tau),$ (12)
@@
$E_{r,i}=\frac{1}{2}\left(\frac{\mathrm{d}x_{i}}{\mathrm{d}\tau}\right)^{2}+U_{\rm eff}(x_{i};\ell_{i},\tau)$ (13)
@@
$\frac{\mathrm{d}E_{r,i}}{\mathrm{d}\tau}=-\Gamma\left(\frac{\mathrm{d}x_{i}}{\mathrm{d}\tau}\right)^{2}+\partial_{\tau}U_{\rm eff}.$ (14)
@@
Alt text: A radial potential curve remains near zero at small dimensionless radius, rises sharply at the shell location $x_{0}$, reaches a narrow maximum labelled as the gap barrier $x_{\rm g}$, and then falls to an outer minimum labelled as the ring-supporting well $x_{\rm r}$.
@@
with log-uniform sampling for $\mu$, $\ell$, and $\Delta$, uniform sampling for $x_{0}$ and $\epsilon$, and $\Delta/x_{0}\leq 0.5$. For each realization all internal critical points of $W$ were located and classified. The full numerical recipe, grid convergence checks, object-level filtering, and the Monte Carlo catalogue format are given in the online supplementary material.
@@
Alt text: Scatter plot of Monte Carlo candidates in the $x_{\rm g}/x_{0}$$x_{\rm r}/x_{0}$ plane. Most blue points lie inside the localization window, concentrated just to the right of $x_{\rm g}/x_{0}=1$, with $x_{\rm r}/x_{0}$ greater than one.
@@
respectively (Schweizer et al., 1987; Finkelman et al., 2011; Finkelman and Brosch, 2011; Pakdil et al., 2017). All three fall within the central model interval. JO171, whose ring has been interpreted as externally formed and is now being reshaped by ram-pressure stripping in a cluster environment, has
@@
The gap and ring definitions, adopted uncertainty bands, Hoag A/B gap choices, and velocity-normalized $j_{\rm phys}$ distributions are tabulated in the online supplementary material.
@@
Alt text: Two-panel comparison of model and observed gap-to-ring ratios. The upper histogram peaks within the shaded central interval from $q=0.677$ to $0.906$ with median $q=0.818$; the lower panel shows Hoag, UGC 4599, and PGC 1000714 overlapping this interval, while JO171 lies at much smaller $q$.
@@ Acknowledgements
We thank the anonymous referee for helpful comments.
@@
The observational quantities used in this work are taken from the cited literature. The numerical catalogues and scripts generated for this study are available in the public GitHub reproducibility repository: https://github.com/Scarlettxy3/hoag-boundary-layering.